OpenStudy (anonymous):

In the center of town, there is a square park with side length of 30 ft. If a person walks from one corner of the park to the opposite corner, how far does the person walk? Round to the nearest foot. 21 42 52 60

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

a^2 + b^2 = c^2 30^2 + 30^2 = x^2 solve for x